A 50 g bracelet is suspected of being gold plated lead instead of pure gold, when it is dropped into a full glass of water, 4.0 cm3 of water overflows. Is the bracelet pure gold? If not, what is the percentage of gold in it? Density of gold is 19 g/cm3  density of lead is 11 g/cm3  



1
Expert's answer
2021-02-15T00:57:08-0500

(a) The volume of water displaced equals the volume of the bracelet. Then, we can find the density of the bracelet:


"\rho=\dfrac{m}{V}=\dfrac{50\ g}{4.0\ cm^3}=12.5\ \dfrac{g}{cm^3}."

The density of gold equals "19\ \dfrac{g}{cm^3}", therefore, the bracelet isn't made of pure gold.

(b) Let the fraction of gold be "x". Then, we can write the following equation:


"19x+11(1-x)=12.5,""8x=1.5,""x=\dfrac{1.5}{8}\times100\%=18.75\%."